OpenStudy (anonymous):

Use the commutative property to write an equivalent numerical expression: 7 + 13 =

OpenStudy (anonymous):

What is the commutative property of addition?

OpenStudy (anonymous):

how do I find that?

OpenStudy (anonymous):

Well its been around fro a very long time. I'm sure wikipedia as a page stating what it is. This is what it is given any two numbers, say a and b, then the commutative property of addition states that a + b = b + a

OpenStudy (anonymous):

so the property of addition is 7 and 13?

OpenStudy (anonymous):

The commutative property for addition is that a + b = b + a. So, for example, 1 + 2 = 2 + 1. Intuitively, the commutative property tells you that it doesn't matter whether you add 1 to 2, or instead, add 2 to 1; they mean the same thing.

OpenStudy (anonymous):

7 + 13 = 13 + 7

OpenStudy (anonymous):

Bingo, that's precisely what they are looking for.

OpenStudy (anonymous):

oh that's it? thank you

OpenStudy (anonymous):

No problem